# WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking

### v4

@@ Abstract
Large Language Models (LLMs) frequently output the label Unknown in reasoning tasks, where two scenarios may appear: (i) an input sample is genuinely unverifiable, but the model cannot understand why; and (ii) a verifiable problem that the model fails to solve, and thus the model outputs Unknown. We refer to these cases collectively as the Vague Perception phenomenon. Current evaluations focus on whether such answers are honest, rather than analyzing the limits of LLM reasoning. To address this, we introduce WakenLLM, a framework that quantifies the portion of Unknown output attributable to model incapacity and evaluates whether stimulation can convert them into either correct answers (verifiable samples) or justified explanations (unverifiable samples) supported by valid reasoning. Our method offers a clearer picture of the limits of LLM reasoning and the potential for corrections across various datasets. Comprehensive experiments on six LLMs suggest that, without further training or parameter revision, LLMs can achieve up to a 68.53% increase in accuracy on Vague Perception samples through guided understanding. Our work reveals that current baseline methods only activate a small portion of LLMs’ reasoning potential, indicating considerable unexplored capacity. This extends the empirical upper bounds of reasoning accuracy in LLMs. Consequently, this study deepens our understanding of the latent reasoning capacity of LLMs and offers a new perspective on addressing the Vague Perception phenomenon.
WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking
Zipeng Ling2∗, Yuehao Tang1∗, Shuliangliu1∗, Junqi Yang3∗, Shenghong Fu5 Chen Huang3, Kejia Huang4, Yao Wan3, Zhichao Hou2, Xuming Hu1 1Hong Kong University of Science and Technology (Guangzhou), 2University of Pennsylvania 3Huazhong University of Science and Technology 4Nanjing University of Posts and Telecommunications, 5Hong Kong Polytechnic University {zpling0816, xuminghu97}@gmail.com
Figure 1: Definition of Vague Perception: when an LLM misjudges a verifiable sample as Unknown, or it is unable to tell why a sample should be categorized as unknown when facing objectively unsolvable ones. The figure presents two representative reasons attributed to this phenomenon: Fact Understanding and Reasoning Gap—highlighting 1. Limited ability to solve a problem. 2. Misunderstandings of facts versus failures to link them.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT‑3, PaLM, and GPT‑4 have shown impressive zero‑shot and few‑shot reasoning on benchmarks like BIG‑Bench, MMLU, and GSM8K Brown et al. (2020); OpenAI (2023); Chowdhery et al. (2022); Srivastava and et al. (2022); Hendrycks and et al. (2021); Cobbe et al. (2021). Their performance is further improved by prompting techniques such as Chain-of-Thought, Self‑Consistency, and Program‑Aided Reasoning Kojima et al. (2022); Wei et al. (2022b); Wang et al. (2023c); Lewkowycz and et al. (2022). Meanwhile, most evaluations rank the overall trustworthiness the first and causal atmosphere secondJiang et al. (2024); Bubeck (2023). Recent diagnostic benchmarks—such as potential‑space reasoning Hagendorff and Fabi (2025), visual hallucination stability Liu et al. (2025), and metaphysical reasoning (MARS) Wang and Song (2024)—demonstrate the need for more fine‑grained evaluations.
Table 1: Comparison of different LLM reasoning–guidance frameworks. WakenLLM integrates all eight dimensions as the first holistic solution for diagnosing and mitigating the Vague Perception phenomenon. For used symbols, : Fully covered, : Partially covered, : Not covered.
Paradigm Guidance Questioning Stimulation
Dynamic Samples
Selection
Reflection
Stage-wise
Examining
Fine-Grained
Decomposing
Diversity
CoT Wei et al. (2022b)
ToT Yao et al. (2023b)
Faithful-CoT Lyu et al. (2023)
Plan&Solve Wang et al. (2023b)
Self-Refine Madaan et al. (2023)
Reflexion Shinn et al. (2023)
Decomposed Prompting Khot et al. (2023)
Active Prompt Diao et al. (2023)
SCORE Zhang et al. (2024)
ProCo Wu et al. (2024)
CoTAL Cohn et al. (2025)
Cochain Zhao et al. (2025)
ISP2 Zhu et al. (2025)
Trace-of-Thought McDonald and Emami (2025)
MAPS de Souza Loureiro et al. (2025)
CoD Xu et al. (2025)
ReSearch Chen et al. (2025)
@@
Fine-grained reasoning reliability remains under-explored. Empirical studies have shown that LLMs frequently hallucinate, or output “Unknown” even when the context contains sufficient evidence, while at other times they output Unknown without understanding why, facing fundamentally under-specified Lin (2023); Zhou et al. (2025) problems. This paper formalizes this phenomenon as Vague Perception: a mismatch between a model’s reasoning ability and the objective problems. Therefore, a critical gap lies in understanding how LLMs deal with complex reasoning and uncertainty.
In this work, we analyze the root causes of deep-reasoning failures behind Vague Perception and propose fine-grained mitigation strategies. To enable systematic measurement, we introduce a new benchmark that spans four subject domains and multiple subcategories, offered in two different formats: (i) Fact-based benchmark, where relevant information is listed as discrete atomic statements (e.g., “A is X, B is Y”); and (ii) Story-based benchmark, where the same facts are embedded in a coherent passage. This design probes how presentation style and information density influence the emergence and eventual resolution of Vague Perception phenomenon.
We construct benchmarks in which half of the samples are verifiable samples with T/F labels and the other half are unverifiable samples with Unknown labels that are objectively unsolvable. These samples are then fed into the model to extract the corresponding Vague Perception samples. As shown in Figure 1, LLMs sometimes output “Unknown” without a clear understanding of the sample. Subsequently, the selected ones are processed via a two-stage stimulation mechanism aimed at “waking up” the model’s reasoning abilities. By measuring “Awaken Ability", we present a novel perspective on evaluating LLM reasoning, one that considers not only the model’s initial accuracy, but also its potential to be guided for correction.
@@
We propose WakenLLM as a strong baseline to explore the reasoning abilities of LLMs among Vague Perception samples, yielding comparable performance across six different LLMs in these reasoning tasks, highlighting that both open-source and closed-source LLMs show substantial improvements in reasoning accuracy.
To summarize, our contributions are threefold:
  • We dive into Vague Perception, a phenomenon in reasoning tasks that models output unknown because of subjective incapacity. To examine it comprehensively, we propose WakenLLM, a unified framework that quantifies LLMs’ capacity to convert to correct answers with stimulation—offering a principled solution for handling Vague Perception in reasoning tasks.
  • Through systematic analyses across LLMs of varying parameter scales, we devise benchmark metrics that more precisely characterize the challenges posed by Vague Perception. The resulting insights not only boost reasoning accuracy but also deepen our understanding of LLM trustworthiness and advanced reasoning paradigms.
@@
  • By comparing WakenLLM with diverse baselines, we measure how effectively each approach unlocks LLMs’ potential. Incorporating both training and non-training frameworks, our work offers a new perspective on mitigating the Vague Perception phenomenon.
Figure 2: Overview of the proposed pipeline. We begin with a balanced Fact and Story-based dataset, sampling from verifiable (T/F) and unverifiable (Unknown) samples at 1 : 1 ratio. The pipeline then (1) detects cases that expose the model’s Vague Perception during reasoning and (2) applies a two-stage remedy: Stage 1 uses stimulation to re-evaluate the mislabeled samples, while Stage 2 gathers the remaining False Converting samples from Stage 1 for model self-reflection. Stimulation Types vary with data form.
2 Evaluation Framework
The WakenLLM evaluation framework, illustrated in Figure 2, is composed of two successive stages. It evaluates the model’s latent reasoning ability and systematically evaluates the challenge of Vague Perception. For the vanilla setting, the first stage assesses the initial Vague Perception samples, while the second investigates the false converting samples $FC^{2}$, allowing us to compute the Overall Benchmark Converting Rate, indicating the percentage of Vague Perception that can be correctly reasoned by LLMs. Under the Remind-then-Guide (RtG) setting, in addition to reminding the model of its prior mistake, we either assign labels—true or false with a preset probability—or provide the previous reasoning process as input to examine whether the LLM still exhibits conformity. We examine whether LLMs still show conformity after being reminded that they solved the problems wrong previously, or it would learn from it and distinguish whether it is guiding or misguiding.
2.1 Overall Benchmark Converting Rate
Our framework begins with a model-specific pre-processing stage that identifies Vague Perception samples—cases where the LLM initially mislabels (e.g., saying Unknown when the sample is verifiable, or failing to explain an unverifiable sample). Since this process is performed separately for each model, the set of Vague Perception samples differs across each model.
After feeding the Stage 1 pipeline with Vague Perception samples, the conversion results fall into three categories: (1) The model fails to generate a coherent logic chain and remains output Unknown, which we call Unexcited Converting (UC). (2) The model produces a false logic chain and an incorrect final label, denoted as False Converting (FC). (3) The model outputs a correct label under stimulation, called True Converting (TC). To quantify these proportions, we use UCR, FCR, and TCR to represent the rates of each type of conversion. Given our focus on the reasoning correction of the model, we elaborate only on TCR and OCR, defining True Converting Rate(TCR) in n-th stage as follows:
$TC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbf{1}\bigl(\hat{y}_{f}^{n}=y_{f}\bigr),\quad TCR_{f}^{n}\;=\;\frac{\bigl|TC_{f}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}\,$
In the above equations, $f\in\{u,v\}$ denotes the sample type (verifiable or unverifiable), and $n\in\{1,2\}$ indicates the stimulation stage. $D_{\text{VP}}^{n}$ represents the input of Vague Perception samples in the $n$-th stage. First stage input $D_{\mathrm{VP}}^{1}$ was extracted from the reasoning dataset. For the second stage, we collect the False Converting part samples as the input, from which we can obtain: $D_{\mathrm{VP}}^{2}=FC^{1}$.
It should be noted that both $TCR_{v}$ and $TCR_{u}$ quantities represent the sum of results from stage 1 and 2, and $TCR^{1}$ denotes the converting rate of both forms data in first stage. To be specific, $TCR_{v}$ = $TCR_{v}^{1}$ + $TCR_{v}^{2}$, and $TCR^{1}$ = $TCR_{v}^{1}$ + $TCR_{u}^{1}$. To integrate performance across both stages, we define Overall Converting Rate (OCR) as follows:
$OCR=\frac{\lvert TC^{1}\cup TC^{2}\rvert}{\lvert D_{\mathrm{vp}}\rvert},\quad TC_{1}\cap TC_{2}=\emptyset$
OCR refers to the portion of under-stimulated samples that turn out to be predicted correctly. By doing this, we can obtain the latent accuracy of LLMs.
2.2 Remind-then-Guide (RtG) Conformity
Label Conformity
We test whether Vague Perception samples display better or worse stability—i.e. whether they are more easily misguided—relative to ordinary settings indicated in Weng et al. (2025). According to previous work, all models demonstrate Conformity. In this part, we first inform the model that it has misunderstood the sample, and then provide a randomly assigned label with varying probabilities, to test whether this misguiding affects its final decision.
We evaluate this phenomenon by setting up one metric for two types of samples: f can be v or u, and we split the results by whether the sample is verifiable or unverifiable. We compute the difference in accuracy between completely wrong and completely correct label guidance and then average the results. Detailed Remind-then-Guide conformity tests on assigning misguiding labels at four given possibilities (100%, 50%, 33.3%, 0%) is available in the Table 891011.
@@ 2.2 Remind-then-Guide (RtG) Conformity
Figure 3: Overall benchmark distribution, which contains 2,840 samples grouped into two forms: story-based (56.3%) and fact-based (43.7%). A detailed breakdown of all topics and “Else” category is provided in the Appendix B.
Reasoning Process (RP) Conformity
We examine how effectively the reasoning chain generated in Stage 1 serves as inspiration or guidance for Stage 2 when it is supplied as additional input. To ensure that the model demonstrates genuine understanding rather than simply being prompted to output the correct answer by accident, we propose two metrics to dive deeper into this phenomenon.
Correct Guiding Rate (CGR) CGR assesses whether the reasoning process output in Stage 1 can assist Stage 2 in stimulating samples to better avoid prior mistakes. Similarly. We remind the LLMs that these explanations may contains flaws because the samples come from the False Converting set. CGR is calculated by the changes of $TCR^{2}$ with and without Reasoning Process.
Reasoning Process Coherence (RPC) This involves re-running Stage 1 with its own prior reasoning process to assess stability and coherence. Similarly, RPC is calculated by changes of $TCR^{1}$ with and without Reasoning Process.
For both metrics, RPC and CGR, below zero indicates that the previous reasoning process is misguiding and results in declined accuracy, while above zero suggests that it is helpful and can guide the LLM towards correct reasoning.
Table 2: Basic info of each dataset.
@@
Table 3: Comparison of six LLMs across four datasets (FLD, FOLIO, Science, Arts), grouped by evaluation settings. RP refers to Reasoning Process under Remind-then-Guide setting. The highest value is marked darker per column.

### v5

# Awakening LLMs’ Reasoning Potential: A Fine-Grained Pipeline to Evaluate and Mitigate Vague Perception

@@ Abstract
Large language models (LLMs) are increasingly trained to abstain on difficult questions by answering unknown. However, we observe that LLMs often misuse this option: they output unknown even when LLMs can actually solve the questions, or they fail to understand why questions are truly unsolvable. We formalize this mismatch between potential ability and the inclination of abstention as the Vague Perception phenomenon. We introduce the WakenLLM pipeline that (1) Extracts Vague Perception samples and (2) Measures how many of them can be converted to correct answers under stimulation. Based on stage-wise metrics (TCR, OCR, etc.) and the upper-bound accuracy Acc(WakenLLM), we quantify LLMs’ reasoning potential beyond one-shot accuracy. Experiments on six LLMs suggest that, without further training or parameter revisions, LLMs can achieve up to a 68.53% increase in accuracy on Vague Perception samples through our designed pipeline. We further analyze how Vague Perception, Conformity and Degradation vary from model families and parameter sizes, and offer model selection strategies in multi-stage reasoning workflows. Finally, by comparing WakenLLM against mainstream reasoning baselines, both training and non-training ones, we show that existing baselines only activate a small portion of LLMs’ reasoning potential, pointing to perception-aware reasoning as a promising direction for future LLM designing. Code and datasets are available at https://github.com/WakenLLMTeam/WakenLLM-toolkit.
Awakening LLMs’ Reasoning Potential: A Fine-Grained Pipeline to Evaluate and Mitigate Vague Perception
Zipeng Ling2∗, Yuehao Tang1∗, Shuliang Liu1∗, Junqi Yang3∗, Shenghong Fu5 Chen Huang3, Kejia Huang4, Yao Wan3, Zhichao Hou2, Xuming Hu1 1The Hong Kong University of Science and Technology (Guangzhou) 2University of Pennsylvania, 3Huazhong University of Science and Technology 4Nanjing University of Posts and Telecommunications 5The Hong Kong Polytechnic University {zpling0816, xuminghu97}@gmail.com
Figure 1: Definition of Vague Perception: LLMs misjudge solvable samples as unknown, or unable to understand unsolvable ones (Left). We attribute this phenomenon to two representative causes: Fact Understanding and Reasoning Gap—highlighting 1. Limited ability to logically solve problems. 2. Limited understanding of facts (Right).
@@ 1 Introduction
LLMs such as GPT‑4, Gemini-2.5 and Claude-3 have shown superior reasoning performance on datasets like BIG‑Bench, MMLU, and GSM8K brown2020language; openai2023gpt4; chowdhery2022palm; srivastava2022beyond; hendrycks2021measuring; cobbe2021training. With stronger performance, many papers focus on the trustworthiness and transparent reasoning to better understand the reasoning paradigmschujie2024honestllm; bubeck2023sparks. Specifically, solving Knowledge Boundary problems can help LLMs to know when to answer and when to abstainli2025knowledgeboundarylargelanguage; Wen2024PerceptionKnowledgeBoundary. However, most prior work focuses on whether and when LLMs should refuse to answer when facing difficult problems. In contrast, we ask a complementary yet opposite question: Are there situations that LLMs give up too early, despite being capable of solving the input questions? If so, inclinations towards abstention would harm LLMs’ ability and performance.
Table 1: Comparison of different LLM reasoning–guidance frameworks. WakenLLM integrates all eight dimensions as the first holistic solution for diagnosing and mitigating the Vague Perception phenomenon. For used symbols, : Fully covered, : Partially covered, : Not covered. Detailed Related Work section can be found in Appendix A.
Paradigm Guidance Questioning Stimulation Dynamic Samples Selection Reflection Stage-wise Examining Fine-Grained Decomposing Diversity
CoT wei2022chain
ToT yao2023tot
Faithful-CoT lyu2023faithful
Plan&Solve wang2023planandsolve
Self-Refine madaan2023selfrefine
Reflexion shinn2023reflexion
Decomposed Prompting khot2023decomposedpromptingmodularapproach
Active Prompt diao2023activeprompt
SCORE zhang2024strongverifier
ProCo wu2024proco
CoTAL cohn2025cotal
Cochain zhao2025cochain
ISP2 zhu2025isp2
Trace-of-Thought mcdonald2025traceot
MAPS loureiro2025maps
CoD xu2025cod
ReSearch chen2025research
@@
We dive deeper into questions that contain uncertainties, as there are two possible situations regarding LLMs outputting unknown: (1) Problems are objectively unsolvable, or (2) LLMs subjectively and wrongly consider the problems unsolvable. The inclination of outputting unknown can largely harm LLMs’ performance, as this situation could be potentially corrected, this paper formalizes it as theVague Perception phenomenon, as in Figure 1. Since potential reasoning ability remain under-explored, understanding how LLMs express uncertainty and the potential of overcoming this phenomenon can be rather important for improving trustworthiness and reliable LLM reasoning.
@@
To address this, we propose WakenLLM as a fine-grained pipeline to explore potential reasoning abilities of LLMs on Vague Perception samples, yielding substantial improvements in accuracy across six LLMs, both open-source and closed-source families, which further proves that LLMs’ reasoning ability is harmed by inclinations of outputting unknown and abstention.
Dataset samples are input into LLMs for inference, and we extract the corresponding Vague Perception samples that meet the conditions in Figure 1. Subsequently, the samples are processed via a two-stage pipeline aiming at “awakening” LLMs’ reasoning potential against the Vague Perception phenomenon, as shown in Figure 3. By measuring the “awakening ability", we present a different perspective on evaluating and mitigating this phenomenon, one that considers not only LLMs one-shot performance, but also the potential ability of solving problems instead of abstaining.
Apart from evaluations and problem mitigation, we offer model-selection strategies in multi-stage reasoning workflows, as our pipeline shows that placing specific models at certain sequences matters much. Besides, we dive deeper into the Vague Perception phenomenon to explore the root causes. To summarize, our contributions are fourfold:
  • We formulate the Vague Perception phenomenon, where models output unknown despite having the potential to solve problems correctly. We propose WakenLLM, a fine-grained pipeline that evaluates the percentage of Vague Perception samples, and mitigate this phenomenon by eliciting LLMs’ reasoning potential.
  • We introduce multiple stage-wise converting metrics (TCR, OCR, etc.) and the overall upper-bound accuracy Acc(WakenLLM) that quantify the ratio of Vague Perception samples can be converted to correct answers without any further training or parameters revisions, evaluating and eliciting LLMs’ reasoning potential.
  • Across six LLMs and four datasets (Fact-based and Story-based) from different fields, we analyze how Vague Perception, Conformity, and Degradation vary from model families and parameter sizes, thus we offer LLM selection strategies for placing models at specific sequences in multi-stage reasoning workflows.
@@
  • By comparing WakenLLM with existing reasoning baselines (both Training and Non-Training), we measure how effectively each framework elicits LLMs’ potential. Our work offers a new perspective on future developing LLM reasoning frameworks with more focus on their under-explored potential.
@@
Figure 2: Overall dataset distribution, with samples grouped into two forms: story-based (56.3%) and fact-based (43.7%). Details of topics (“Else” category included) is in Appendix C.
2 Datasets Composition
To explore the effects of the Vague Perception phenomenon on different types of samples, we classify them by their labels: (1) solvable samples, where problems have True/False labels that can be solved with valid reasoning; and (2) unsolvable samples, which are labeled unknown and objectively can not be solved. Furthermore, as in Figure 2, to better understand how different context styles influence, we categorize datasets as follows:
Fact-based Datasets
The information in contexts is stated atomically, like “A is X, B is Y". FLD (morishita2024enhancing) supplies three kinds of sample labels annotated as True, False, or Unknown. We randomly sample 600 samples: 300 solvable (True/False) and 300 unsolvable Unknown ones. And we draw 640 samples—320 solvable samples with T/F labels and 320 Unknown samples from FOLIO (han2022folio), which has the same three-way labeling.
Story-based Datasets
The contexts are long passages, more aligned with humans’ reading habits. We extract samples from four fields: Physics, Biology, Figurative-language and Writing-strategies from ScienceQA (lu2022learn), and applied the same sampling strategy. The former two are categorized as Science datasets, the latter two are Arts datasets. Dataset details are in Appendix C.
Table 2: Information of datasets used in experiments.
@@ 2 Datasets Composition
Figure 3: Overview of our proposed pipeline. We begin with Fact and Story-based datasets, sampling from solvable (T/F) and unsolvable (Unknown) samples at 1 : 1 ratio. The pipeline then (1) Extracts Vague Perception samples during LLMs’ reasoning and (2) Applies a two-stage pipeline: stage 1 uses stimulation to re-evaluate the mislabeled samples, while stage 2 gathers the remaining False Converting samples from stage 1 for model self-reflection. Stimulation types vary from different dataset forms.
3 Proposed Pipeline: WakenLLM
The WakenLLM pipeline is illustrated in Figure 3. It first extracts Vague Perception samples from input datasets, later the two-stage stimulation is implemented to elicit model’s potential on these samples. For the Vanilla Setting, the stage 1 deals with the initial Vague Perception samples, and the stage 2 receives falsely answered samples from stage 1 ($FC^{1}$). Via the two-stage stimulation pipeline, we can get the percentage of Vague Perception samples that are eventually corrected by LLMs. For Remind-then-Guide (RtG) settings, in addition to reminding LLMs of prior mistakes, we either randomly assign labels, or provide previous reasoning processes in the prompt. We examine whether LLMs still show conformity after being reminded of answering problems wrong previously; or they would conform to previous reasoning processes and make the same mistakes. Detailed metrics and pseudo-code visualizations are available in Appendix D and M.
3.1 Samples Extraction
Our pipeline begins with reasoning on pre-processed datasets with 50% solvable (Label:True/False) and 50% unsolvable (Label:Unknown) samples, we acquire tasks accuracy and extract corresponding Vague Perception samples—where LLMs output unknown when samples are solvable, or fail to explain why samples are truly unsolvable. Since this is performed separately for each model, Vague Perception samples differs from each too. Details are in Appendix J and Table 8.
3.2 Vanilla Setting
After extracting Vague Perception samples, we input them into stage 1. The conversion results fall into three categories: (1) LLMs output correct labels under stimulation, called True Converting (TC). (2) LLMs output an incorrect T/F label for solvable samples, denoted as False Converting (FC). (3) LLMs fail to comprehend and remain output Unknown for solvable samples, which is Unknown Converting (UC). To quantify these proportions, we use UCR, FCR, and TCR to represent rates of each type converting. Detailed formulas are in Appendix F. Given our focus on the reasoning correction of LLMs, we define them as:
$TC_{t}^{n}\;=\;\sum_{y_{t}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbbm{1}\bigl(\hat{y}_{t}^{n}=y_{t}\bigr),\quad TCR_{t}^{n}\;=\;\frac{\bigl|TC_{t}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}\,$
$FC_{t}^{n}\;=\;\sum_{y_{t}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbbm{1}\bigl(\hat{y}_{t}^{n}\neq y_{t})$
$UC_{t}^{n}\;=\;\sum_{y_{t}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbbm{1}\bigl(\hat{y}_{t}^{n}\in U,\;y_{t}\notin U\bigr)$
In the above equations, $t\in\{s,u\}$ denotes sample types (solvable or unsolvable), and $n\in\{1,2\}$ indicates the stimulation stage. $D_{\text{VP}}^{n}$ represents the input of Vague Perception samples in the $n$-th stage. First stage input $D_{\mathrm{VP}}^{1}$ was extracted from the input datasets. For the second stage, we collect the False Converting samples as the input, from which we can obtain: $D_{\mathrm{VP}}^{2}=FC^{1}$.
Omitting superscripts like $TCR_{s}$ represent the sum of results from stage 1 and 2, and $TCR^{1}$ denotes the converting rate of both sample forms in the first stage. To be specific, $TCR_{s}$ = $TCR_{s}^{1}$ + $TCR_{s}^{2}$, standing for the TCR value of solvable samples in both stages; Likewise, $TCR^{1}$ = $TCR_{s}^{1}$ + $TCR_{u}^{1}$, indicating TCR for stage 1 that sums up both solvable and unsolvable samples.
To get a comprehensive view of LLMs’ reasoning potential, we define Overall Converting Rate (OCR) as follows:
$OCR=\frac{\lvert TC^{1}\cup TC^{2}\rvert}{\lvert D_{\mathrm{vp}}\rvert},\quad TC^{1}\cap TC^{2}=\emptyset$
OCR refers to the portion of Vague Perception samples that are eventually corrected by LLMs. By calculating this, we can obtain the potential accuracy harmed by inclinations of abstention.
3.3 Remind-then-Guide (RtG) Setting
Label Conformity
We test how LLMs display conformity facing Vague Perception samples. We firstly inform LLMs of inputting Vague Perception samples, then provide randomly assigned labels to test whether misguiding affects outputs. We compare TCR changes with and without the RtG setting, and average results on two stages:
$Conf_{t}=\frac{1}{2}(\underbrace{TCR^{1}_{t}-TCR_{t}^{1^{\prime}}}_{\text{Stage 1 Conformity}}+\underbrace{TCR^{2}_{t}-TCR_{t}^{2^{\prime}}}_{\text{Stage 2 Conformity}})$
Table 3: Comparison of six LLMs across four datasets (FLD, FOLIO, Science, Arts), grouped by evaluation settings. RP refers to reasoning process under Remind-then-Guide setting. The highest value is in bold per column.